ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 207, 2015

Aizenbud, A. and V. Drinfeld The wave front set of the Fourier transform of algebraic measures

527

Andrews, U., M. Cai, D. Diamondstone, S. Lempp and J. S. Miller On the structure of the degrees of relative provability

449

Araújo, J. a., W. Bentz and J. Konieczny The commuting graph of the symmetric inverse semigroup

103

Avila, A., D. Dolgopyat, E. Duryev and O. Sarig The visits to zero of a random walk driven by an irrational rotation

653

Baek, S. On the torsion of Chow groups of Severi–Brauer varieties  899

Bentz, W. See J. a. Araújo Bogart, T., C. Haase, M. Hering, B. Lorenz, B. Nill, A. Paf-

fenholz, G. Rote, F. Santos and H. Schenck Finitely many smooth d-polytopes with n lattice points

301

Bourgain, J. and Z. Rudnick Nodal intersections and L^{p} restriction theorems on the torus

479

Bowen, L., R. Grigorchuk and R. Kravchenko Invariant random subgroups of lamplighter groups

763

Cai, M. See U. Andrews Can, M. B., Y. Cherniavsky and T. Twelbeck Lexicographic

shellability of the Bruhat–Chevalley order on fixed-point-free involutions

281

Cherniavsky, Y. See M. B. Can Crestani, E. and A. Lucchini Bias of group generators in the solv-

able case

739

D’Alessandro, S., P. Hájek and M. Johanis Corrigendum to the paper “Polynomial algebras on classical Banach spaces”, Israel Journal of Mathematics 106 (1998), 209–220

1003

Diamondstone, D. See U. Andrews

Dirbák, M. Minimal extensions of flows with amenable acting groups  581

Dolgopyat, D. See A. Avila Drinfeld, V. See A. Aizenbud Duryev, E. See A. Avila

Dydak, J. and M. Levin Extensions of maps to Moore spaces  981
Eldan, R. An efficiency upper bound for inverse covariance estimation  1

Elduque, A. and M. Kochetov Graded modules over classical simple Lie algebras with a grading

229

Ginzburg, D. and J. Hundley A doubling integral for G_{2}  835

Grigorchuk, R. See L. Bowen Haase, C. See T. Bogart Hagen, M. F. and P. Przytycki Cocompactly cubulated graph

manifolds

377

Hájek, P. See S. D’Alessandro Hamkins, J. D., G. Leibman and B. Löwe Structural connections

between a forcing class and its modal logic

617

Hering, M. See T. Bogart Hornbostel, J. and S. Schröer Infinite CW-complexes, Brauer

groups and phantom cohomology

77

Huang, L. Ricci curvatures of left invariant Finsler metrics on Lie groups

783

Hundley, J. See D. Ginzburg Johanis, M. See S. D’Alessandro

Kemarsky, A. Distinguished representations of GL_{n}(ℂ)  435

Khukhro, E. I. and P. Shumyatsky Nonsoluble and non-p-soluble length of finite groups

507

Kida, Y. Stable actions of central extensions and relative property (T)  925

Kravchenko, R. See L. Bowen Kochetov, M. See A. Elduque Konieczny, J. See J. a. Araújo La Mattina, D. Almost polynomial growth: classifying varieties of

graded algebras

53

Leibman, G. See J. D. Hamkins Lempp, S. See U. Andrews Levin, M. See J. Dydak Lorenz, B. See T. Bogart Löwe, B. See J. D. Hamkins Lucchini, A. See E. Crestani Marcus, B. and R. Pavlov An integral representation for topolog-

ical pressure in terms of conditional probabilities

395

Massart, D. Erratum to “On Aubry sets and Mather’s action functional”, Israel Journal of Mathematics 134 (2003), 157–171

1001

Matsuo, S. and M. Tsukamoto Local mean dimension of ASD moduli spaces over the cylinder

793

Miller, J. S. See U. Andrews

Moriya, K. A factorization of a super-conformal map  331

Morris, I. D. A note on configurations in sets of positive density which occur at all large scales

719

Nikolayevsky, Y. Totally geodesic hypersurfaces of homogeneous spaces

361

Nill, B. See T. Bogart Paffenholz, A. See T. Bogart Pavlov, R. See B. Marcus

Procesi, C. On the theorem of Amitsur–Levitzki  151

Przytycki, P. See M. F. Hagen Rabier, P. J. Differentiability of quasiconvex functions on separable

Banach spaces

11

Rote, G. See T. Bogart Rudnick, Z. See J. Bourgain Santos, F. See T. Bogart Sarig, O. See A. Avila Schenck, H. See T. Bogart Schröer, S. See J. Hornbostel

Shimizu, K. On indicators of Hopf algebras  155

Shumyatsky, P. See E. I. Khukhro Skryabin, S. Invariant subrings and Jacobson radicals of Noetherian

Hopf module algebras

881

Tao, T. and T. Ziegler A multi-dimensional Szemerédi theorem for the primes via a correspondence principle

203

Tsukamoto, M. See S. Matsuo Twelbeck, T. See M. B. Can

Zhang, S. Coarse quotient mappings between metric spaces  961

Ziegler, T. See T. Tao